# Sociodynamics-inspired Adaptive Coalition and Client Selection in Federated Learning

### v1

@@ 4 Experiments: FedCVR-Bolt in Heterogeneous Scenarios
Let us present experimental results evaluating FedCVR-Bolt in heterogeneous FL scenarios. We utilize established FL benchmark datasets caldas2018leaf ; li2020federated and also introduce a synthetic dataset specifically designed for a controlled federated linear regression analysis. A detailed description of all dataset settings is provided in Appendix B. In our analysis, we compare FedCVR-Bolt with several prominent FL client selection algorithms: the uniform random selection policy employed by FedAvg mcmahan2017communication , the exploitative Power-of-Choice strategy cho2022towards which selects clients with higher local losses, and ActiveFL goetz2019active , which probabilistically favors clients with higher losses to foster exploration. Section 4.1 details the synthetic linear regression experiments, configured to demonstrate the effectiveness of the FedCVR-Bolt policy under challenging heterogeneous conditions. Subsequently, Section 4.2 reports classification results on more complex, real-world heterogeneous datasets.
@@ 4.1 Federated Linear Regression
Cluster parameters, client assignments, and train/test splits follow uniform sampling routines, detailed in Appendix B. Table 1 reports the average test MSE after $T=100$ rounds, selecting $P=10$ clients per round with $S=10$ local SGD steps. In IID settings ($J=1$), all policies perform similarly, with FedCVR-Bolt maintaining competitive results. In non-IID scenarios ($J=2$), FedCVR-Bolt consistently outperforms baselines. For instance, it reduces test MSE by up to $3.0\%$ over uniform sampling in the no-intercept, $D=1$ case, and by $1.7\%$ when an intercept is included.
Table 1: Comparison between our algorithm, FedCVR-Bolt , and FL selection baselines on a synthetic regression dataset. We evaluate two cases: $D=1$ (linear model without intercept) and $D=2$ (linear model with intercept), under both IID and non-IID settings. We report the Test MSE, averaged across clients (the lower the better). In the IID case, all methods perform comparably. In the non-IID case, FedCVR-Bolt achieves the lowest MSE, improving performance by approximately 3% for $D=1$ and 1.7% for $D=2$.
Regression Model Heterogeneity FedAvg Power-of-Choice AFL FedCVR-Bolt $y=\theta_{1}x$ IID 7.3621 7.3712 7.3767 7.3710 Non-IID 73.4670 73.3034 73.1661 71.3769 $y=\theta_{0}+\theta_{1}x$ IID 0.0587 0.0584 0.0584 0.0580 Non-IID 2.0554 2.0403 2.1234 2.0195
@@ 4.2 FL Classification Benchmarks
We evaluate FedCVR-Bolt on standard FL benchmarks. The Synthetic(1,1) dataset li2020federated simulates a heterogeneous and non-IID setting with $K=100$ clients, among which $P=10$ are sampled per round for $T=100$ communication rounds, each performing $S=10$ local epochs. We use a logistic regression model for this task. To assess performance in more realistic settings, we conduct experiments on MNIST lecun1998mnist and CIFAR-10 krizhevsky2009learning , using a Dirichlet partitioning with concentration parameter $\alpha=0.1$ caldas2018leaf , introducing label and quantity heterogeneity across clients. Model architectures and hyper-parameters are detailed in Appendix B. Table 2 reports the average test accuracy over all clients hossin2015review . FedCVR-Bolt consistently outperforms the baselines. In particular, it achieves substantial improvements over the random selection policy (FedAvg). The Power-of-Choice strategy, despite being loss-aware, does not yield consistent gains and can even degrade performance in some settings. ActiveFL achieves competitive results but remains below FedCVR-Bolt . Overall, FedCVR-Bolt emerges as the most effective strategy across all settings.
Table 2: Comparison with FL selection baselines on three heterogeneous classification datasets, reporting the test accuracy hossin2015review , averaged across clients (the larger the better): Synthetic(1,1) li2020federated , Federated MNIST lecun1998mnist with $\alpha=0.1$, and CIFAR-10 krizhevsky2009learning with $\alpha=0.1$. FedCVR-Bolt consistently outperforms other client selection methods, achieving higher test accuracy across all benchmarks.
Dataset FedAvg Power-of-Choice Active FL FedCVR-Bolt
Synthetic (1,1) 71.35 76.45 77.00 78.55
MNIST 86.30 81.00 88.89 90.23
CIFAR-10 52.67 49.85 55.19 57.06
@@ 5 Conclusions
Limitations. A key consideration is the computational overhead of clustering client models server-side, particularly in large-scale deployments. Computing client similarity and performing spectral clustering can introduce significant cost, posing scalability challenges. However, these operations are infrequent (e.g., every few rounds), and the cost is often justified by substantial improvements in convergence speed and robustness under heterogeneity.
Broader Impact. FedCVR-Bolt improves FL performance under heterogeneity, benefiting privacy-sensitive areas like healthcare and enabling model tailoring in agriculture. Improved convergence and reduced communication can enhance efficiency in IoT systems. However, identifying and potentially amplifying differences between client groups could inadvertently reinforce biases or be misused for unfair allocation. Responsible deployment requires careful consideration of these risks.
Acknowledgments
A.L. and L.R. worked under the auspices of Italian National Group of Mathematical Physics (GNFM) of INdAM. A.L. was supported by the Project Piano Nazionale di Ripresa e Resilienza - Next Generation EU (PNRR-NGEU) from Italian Ministry of University and Research (MUR) under Grant DM 117/2023. R.R. was supported by the “Higher-order interactions in social dynamics with application to monetary networks" project —funded by European Union— Next Generation EU within the PRIN 2022 program (D.D. 104 — 02/02/2022 The Ministry of University and Research). The authors gratefully acknowledge Politecnico di Torino IT High Performance Computing Resources in the Department of Applied Mathematics, services, and staff expertise.
Appendix A Theoretical Guarantees: Proofs and Details
@@
This appendix provides a comprehensive overview of the theoretical framework behind FedCVR-Bolt . In Section A.1 we introduce the analogy between social dynamics and FL, focusing on the construction of the influence matrix $W$ between clients’ models. In Section A.2, we detail the theoretical foundation of our algorithm, providing formalism and rigor to our design choices. Finally, in Section A.3 we motivate the update policy for each client’s mean given the sampled and observed clients.
A.1 Social Dynamics Setting
@@
In the last decades, the analysis of Social Dynamics has gained significant attention in the scientific community, providing tools to understand, predict, and influence the spread and dissemination of information across a complex network as well as collective decision-making processes. The application of Opinion Dynamics goes beyond this specific application spacing among marketing, politics, decision-making strategies [55]. In order to face classical challenges from Federated Learning, we decide to interpret them through the lens of Social Dynamics. Such an innovative angle allows us to establish new results and gain novel insights on the topic.
@@
In the case of interest, we may interpret each client $k$ as a network node while the node status refers to the client local model $\theta_{k}(t)$ with $t$ index of the communication round. To naturally model the influence matrix $W(t)$ while accounting for the homophily effect, we build upon the approach introduced in [18]. Specifically, we define the social influence of agent $j$ on agent $k$ as
@@
A.2 Sampling Strategy
Once the clusters have been identified, our goal is to determine a representative set of clients to sample, one for each cluster. The approach we adopt is a generalization of classical methods originally developed for sensor placement [46], and later extended to the context of Opinion Dynamics in [19].
@@ A.2 Sampling Strategy
Consider $\theta^{d}\in\mathbb{R}^{K}$ the zero mean vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Given $\mathcal{A}\subseteq K$, it holds [^1] Here, with some abuse of notation, $\mathbb{E}[\theta_{gl}^{d}|\theta_{\mathcal{A}}^{d}]$ denotes the projection of $\theta_{gl}^{d}$ onto the space of all linear combinations of the components of $\theta_{\mathcal{A}}^{d}$ with constant coefficients, found via linear regression. Under standard assumption of normal distribution it would coincide with the true conditional expectation [56].
@@
Recalling federated averaging aggregation [1], at each round the global model is updated as a linear convex combination of local models (refer to Eq.(2)). Furthermore, from classical statistical literature (Section 2.3.1 in [56]) it is well known that the linear least squares projection of the global model $\theta_{gl}^{d}$ given the sampling $\theta_{\mathcal{A}}^{d}$ is computed as
@@
Using the fact that by definition in (2) $\theta_{gl}^{d}=\alpha^{\top}\theta^{d}$, we now derive the following series of equalities:
@@
Thus, based on $v_{\mathcal{A}}^{d}$ definition from (12), applying Lemma 1, defined $\hat{\alpha}$ such that
@@
Despite this metrics allows us to find the best subset of models to sample in order to best predict the global model $\theta_{gl}$, from [19] it is known to be a combinatorial problem which becomes unfeasible increasing the number of samples.
@@ A.3 Update Policy
Appendix B Implementation Details and Further Experiments

### v2

@@ 5 Experiments: FedCVR-Bolt in Heterogeneous Scenarios
We present experimental results evaluating FedCVR-Bolt in heterogeneous FL scenarios. We employ established FL benchmark datasets (Caldas et al., 2018; Li et al., 2020), together with a synthetic dataset specifically designed for a controlled federated linear regression analysis. A detailed description of dataset settings and model architectures is provided in Appendix C.
In our analysis, we compare FedCVR-Bolt against a diverse set of FL baselines. From client selection methods, we include the uniform random policy of FedAvg (McMahan et al., 2017), the exploitative Power-of-Choice strategy (Cho et al., 2022) that favors clients with higher local losses, and ActiveFL (Goetz et al., 2019), which balances exploration and exploitation through probabilistic selection. From the broader FL literature, we also consider FedProx (Li et al., 2020), a regularization-based approach designed for heterogeneous settings, and two clustering-based personalized methods, IFCA (Ghosh et al., 2020) and FeSEM (Long et al., 2023). Section 5.1 reports regression results, highlighting the behavior of selection policies in controlled heterogeneous environments. Section 5.2 then turns to classification tasks on real-world datasets.
@@ 5.1 Federated Linear Regression
Cluster parameters, client assignments, and train/test splits follow uniform sampling routines, detailed in Appendix C. Table 1 reports the average test MSE after $T=100$ rounds, selecting $P=10$ clients per round with $S=10$ local SGD steps. In IID settings ($J=1$), all policies perform similarly, with FedCVR-Bolt maintaining competitive results. In non-IID scenarios ($J=2$), FedCVR-Bolt consistently outperforms baselines. For instance, it reduces test MSE by up to $3.0\%$ over uniform sampling in the no-intercept, $D=1$ case, and by $1.7\%$ when an intercept is included.
Table 1: Comparison between FedCVR-Bolt and FL selection baselines on a synthetic regression dataset. We evaluate two cases: $D=1$ (linear model without intercept) and $D=2$ (linear model with intercept), under both IID and non-IID settings. We report the Test MSE, averaged across clients (the lower the better). In the IID case, all methods perform comparably. In the non-IID case, FedCVR-Bolt achieves the lowest MSE, improving performance by approximately 3% for $D=1$ and 1.7% for $D=2$.
[…]
@@ 5.2 FL Classification Benchmarks
We next evaluate FedCVR-Bolt on standard FL classification benchmarks. We consider MNIST (LeCun, 1998), CIFAR-10, and CIFAR-100 (Krizhevsky, 2009), each partitioned across clients using a Dirichlet distribution with concentration parameter $\alpha=0.1$ (Caldas et al., 2018), which introduces both label and quantity heterogeneity. Each dataset is trained for $T=100$ communication rounds with $P=10$ clients sampled per round and $S=10$ local epochs. Model architectures and hyper-parameters are detailed in Appendix C.
Table 2 reports average test accuracy across clients. FedCVR-Bolt consistently achieves the highest performance across all benchmarks. In particular, it delivers notable gains over random client sampling (FedAvg) and over the loss-aware Power-of-Choice, which can fail to generalize well in heterogeneous regimes. Regularization with FedProx and clustering-based methods such as IFCA and FeSEM yield competitive results, but remain below FedCVR-Bolt . This confirms the effectiveness of our selection strategy in both simple and highly heterogeneous datasets.
Table 2: Comparison with FL baselines on heterogeneous classification benchmarks. We report the test accuracy (Hossin & Sulaiman, 2015), averaged across clients (the larger the better). FedCVR-Bolt consistently outperforms existing methods across MNIST, CIFAR-10, and CIFAR-100, all partitioned with Dirichlet $\alpha=0.1$.
Dataset FedAvg Power-of-Choice Active FL FedProx IFCA FeSEM FedCVR-Bolt MNIST 86.30 $\pm$ 1.12 81.00 $\pm$ 2.10 88.89 $\pm$ 0.98 89.69 $\pm$ 0.91 89.46 $\pm$ 1.04 86.77 $\pm$ 1.85 90.23 $\pm$ 0.82 CIFAR-10 52.67 $\pm$ 1.25 49.85 $\pm$ 1.71 55.19 $\pm$ 1.34 46.44 $\pm$ 1.88 52.94 $\pm$ 1.09 44.66 $\pm$ 2.06 57.06 $\pm$ 0.77 CIFAR-100 22.92 $\pm$ 0.95 23.46 $\pm$ 1.05 23.35 $\pm$ 1.12 22.13 $\pm$ 1.34 22.55 $\pm$ 1.28 11.87 $\pm$ 0.77 24.82 $\pm$ 0.88
@@ 6 Conclusions
7 Reproducibility Statement
The reproducibility of the results and the theoretical contributions of this work has been a paramount concern throughout the entire development of this project and while drafting this manuscript. In this section, we provide details and a concise guide to reproduce our results and verify our contributions.
  • Code Availability: All the code used in our experiments has been included in the supplementary material of the submission. Additionally, we will release online a well-documented and structured final version of the code, to allow for easy reproduction of the experiments detailed in the paper.
  • Datasets and data splits: The datasets used in our experiments are publicly available and can be downloaded online. Detailed instructions on how to access and preprocess the datasets will be provided together with the final code release. The data splits can be generated directly through our provided code. This ensures that others can replicate the exact federated learning scenarios used in our work.
  • Theoretical Results: We provide complete proofs for all theorems and propositions presented in the paper in Appendix B.
We are confident that with these resources, all experimental and theoretical results can be reproduced by the community.
Appendix A Notation
For readability, we summarize below the notation conventions adopted throughout the paper.
Symbol Description
$K$ number of clients
$\mathcal{K}$ set of clients
$\mathcal{A}\subseteq\mathcal{K}$ subset of clients
$W(t)$ influence matrix among clients at round $t$
$T$ number of communication rounds
$D$ number of model parameters
$S$ number of training iterations
$P$ number of participating clients
$\eta$ learning rate
$\theta_{\mathrm{gl}}\in\mathbb{R}^{D}$ global model
$\theta_{\mathrm{gl}}(t)\in\mathbb{R}^{D}$ global model at round $t$
$\theta_{k}(t)\in\mathbb{R}^{D}$ local model of client $k$ at round $t$
$\tilde{\theta}_{k}(t)\in\mathbb{R}^{D}$ normalized local model of client $k$ at round $t$
$\mathcal{L}(\theta)$ global loss
$\mathcal{L}_{k}(\theta)$ client loss
$\mathcal{P}_{t}$ subset of participating clients at round $t$
$\theta_{\mathrm{gl}}^{d}\in\mathbb{R}^{K}$ random vector of $d$-th components of the global model
$\theta^{d}\in\mathbb{R}^{K}$ random vector of $d$-th components of the model
$\bar{\theta}^{d}$ mean of random vector $\theta^{d}$
$\bar{\theta}_{k}(t+1)$ mean of unobserved clients $k\in[K]\backslash\mathcal{C}_{i}(t)$
$C^{d}$ covariance of random vector $\theta^{d}$
$\rho_{jk}$ Pearson correlation coefficient between random variables $j$ and $k$
$v^{d}\in\mathbb{R}^{K}$ value vector for the $d$-th component; $v_{k}^{d}$ is the variance reduction for estimating $\theta_{\mathrm{gl}}^{d}$ when sampling $\theta_{k}^{d}$
$v_{k}$ total variance reduction associated with client $k$
$\Pi(K)$ set of all partitions of $[K]$
$\mathcal{C}_{i}(t)\in\Pi(K)$ $i$-th coalition at round $t$
$\pi_{p}(k;t)$ Boltzmann-like probability that client $k$ is selected in coalition $\mathcal{C}_{i}(t)$
$n_{k}$ number of training samples of $k$-th client
Table 3: Paper notation summary.
Appendix B Theoretical Guarantees: Proofs and Details
@@
This appendix provides a comprehensive overview of the theoretical framework behind FedCVR-Bolt . In Section B.1 we introduce the analogy between social dynamics and FL, focusing on the construction of the influence matrix $W$ between clients’ models. In Section B.2, we detail the theoretical foundation of our algorithm, providing formalism and rigor to our design choices. Finally, in Section B.3 we motivate the update policy for each client’s mean given the sampled and observed clients.
B.1 Social Dynamics Setting
@@
In the last decades, the analysis of Social Dynamics has gained significant attention in the scientific community, providing tools to understand, predict, and influence the spread and dissemination of information across a complex network as well as collective decision-making processes. The application of Opinion Dynamics goes beyond this specific application spacing among marketing, politics, decision-making strategies (Zha et al., 2021). In order to face classical challenges from Federated Learning, we decide to interpret them through the lens of Social Dynamics. Such an innovative angle allows us to establish new results and gain novel insights on the topic.
@@
In the case of interest, we may interpret each client $k$ as a network node while the node status refers to the client local model $\theta_{k}(t)$ with $t$ index of the communication round. To naturally model the influence matrix $W(t)$ while accounting for the homophily effect, we build upon the approach introduced in Mäs et al. (2010). Specifically, we define the social influence of agent $j$ on agent $k$ as
@@
To validate our choice for the proximity matrix, we conduct an ablation study to analyze our framework’s robustness. We implement our algorithm with other similarity functions used in coalition formation (Scholkopf & Smola, 2001):
[…]
  • Sigmoid Kernel: $W_{kj}=\tanh(\gamma\langle\theta_{k}(t),\theta_{j}(t)\rangle+c)$
While our theoretically-motivated choice yields the best performance ($90.23\%$ on MNIST, $57.06\%$ on CIFAR-10), our framework continues to substantially outperform the baselines even with these alternative similarity metrics.
Table 4: Test accuracy (%) across proxy metrics. Best per dataset in bold.
Proxy Metric RBF (ours) Cosine Similarity Laplacian Kernel Sigmoid Kernel
MNIST 90.23% 88.72% 88.28% 88.92%
CIFAR-10 57.06% 55.32% 49.85% 50.25%
B.2 Sampling Strategy
Once the clusters have been identified, our goal is to determine a representative set of clients to sample, one for each cluster. The approach we adopt is a generalization of classical methods originally developed for sensor placement (Das & Kempe, 2008), and later extended to the context of opinion dynamics in Raineri et al. (2025).
@@ B.2 Sampling Strategy
Consider $\theta^{d}\in\mathbb{R}^{K}$ the zero mean vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Given $\mathcal{A}\subseteq K$, it holds
@@
Recalling federated averaging aggregation (McMahan et al., 2017), at each round the global model is updated as a linear convex combination of local models (refer to Eq.equation 2). Furthermore, from classical statistical literature (Section 2.3.1 in Hastie et al. (2009)) it is well known that the linear least squares projection of the global model $\theta_{gl}^{d}$ given the sampling $\theta_{\mathcal{A}}^{d}$ is computed as
@@
Using the fact that by definition in equation 2 it holds $\theta_{gl}^{d}=\alpha^{\top}\theta^{d}$, we now derive the following series of equalities:
@@
Thus, based on $v_{\mathcal{A}}^{d}$ definition from equation 12, applying Lemma 1, defined $\hat{\alpha}$ such that
@@
Despite this metrics allows us to find the best subset of models to sample in order to best predict the global model $\theta_{gl}$, from Raineri et al. (2025) it is known to be a combinatorial problem which becomes unfeasible increasing the number of samples.
@@ B.3 Update Policy
B.4 Convergence Analysis
In this appendix, we establish convergence guarantees for FedCVR-Bolt . Our analysis builds on standard assumptions in FL – smoothness of the global loss and bounded variance of the stochastic federated gradient – extended with a mild alignment condition on the client selection policy. The latter ensures that, in expectation, the update direction preserves a positive correlation with the true gradient, thereby maintaining descent. Under these assumptions, we show that the sequence of global iterates produced by FedCVR-Bolt converges, in expectation, to a neighborhood of a stationary point of the global loss. Here, convergence is meant in the stochastic optimization sense: once averaged over both the randomness of client sampling and the evolution in time, the gradient norm becomes small. This differs from the stronger notion of convergence of the iterates $\theta_{gl}$ themselves to an exact stationary point, which is precluded by the intrinsic variance of stochastic updates but replaced by concentration around regions of vanishing gradient.
Assumption A2.
The global loss function $\mathcal{L}(\theta)$ is $L$-smooth, i.e., it is differentiable with $L$-Lipschitz gradient and its federated gradient $g(t):=\sum_{k\in P_{t}}\tilde{\alpha}_{k}\nabla\mathcal{L}_{k}(\theta_{gl}(t))$ has bounded variance, i.e., $\mathbb{E}_{\mathcal{P}_{t}\sim\pi(t)}[\|g(t)-\mathbb{E}[g(t)]\|^{2}]<\sigma^{2}$.
Assumption A3.
Gradient Variance Alignment. Let $G(\theta):=\mathbb{E}_{\mathcal{P}_{t}\sim\pi(t)}[g(t)]$. Assume there exists a constant $c>0$ s.t. $\langle G(\theta),\nabla\mathcal{L}(\theta)\rangle\geq c\|\nabla\mathcal{L}(\theta)\|^{2}$, i.e., the bias of the policy is informed and points in a descent direction (very weak, and can be verified by our experimental results).
Proposition A3.
Let $\{\theta_{gl}(t)\}_{t}$ the sequence of global model update produced by FedCVR-Bolt algorithm, with known model covariance $C$. Under Assumptions 1 and 2, FedCVR-Bolt, for a small learning rate $\eta>0$, converges to a neighborhood of a stationary point, i.e.,
$\lim_{T\to\infty}\dfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{gl})\|^{2}]<\varepsilon\,.$
Proof.
We focus on a single SGD step, having $\theta_{gl}(t+1)=\theta_{gl}(t)-\eta g(t)$, where $g(t)$ is the federated gradient produced by FedCVR-Bolt. Due to Assumption 1, if the loss is $L$-smooth, we can apply the Descent Lemma, from Chapter 2, Theorem 2.1.5 in Nesterov (2014). Thus, we bound as follows:
$\mathcal{L}(\theta_{gl}(t+1))\leq\mathcal{L}(\theta_{gl}(t))-\eta\langle\nabla\mathcal{L}(\theta_{gl}(t)),g(t)\rangle+\dfrac{L\eta^{2}}{2\|g(t)\|^{2}}.$
By taking the expectation up to the $t$-th round, denoted by $\mathbb{E}_{t}$, our bound becomes:
$\mathbb{E}_{t}[\mathcal{L}(\theta_{gl}(t+1))]\leq\mathcal{L}(\theta_{gl}(t))-\eta\langle\nabla\mathcal{L}(\theta_{gl}(t)),\mathbb{E}_{t}[g(t)]\rangle+\dfrac{L\eta^{2}}{2\mathbb{E}_{t}[\|g(t)\|^{2}]}$
Without losing generality, we can assume that after $T_{0}$ we are in a thermalized regime, where the spectral clustering has converged, and the clusters are stationary, namely $\forall t>T_{0},\,\mathcal{C}_{p}(t)=\mathcal{C}_{p}^{*}$, hence also the Boltzmann measure $\pi(t)$ is stationary. Let us recall that we defined $\mathbb{E}_{t}[g(t)]=G(\theta_{gl}(t))$. Applying Assumption 2, we establish the following bounds:
$\langle\nabla\mathcal{L}(\theta_{gl}(t)),\mathbb{E}_{t}[g(t)]\rangle=\langle\nabla\mathcal{L}(\theta_{gl}(t)),G(\theta_{gl}(t))\rangle\geq c\|\nabla\mathcal{L}(\theta_{gl}(t))\|^{2},$
and
$\mathbb{E}_{t}[\|g(t)\|^{2}]\leq\|G(\theta_{gl}(t))\|^{2}+\mathbb{E}_{t}[\|g(t)-G(\theta_{gl}(t))\|^{2}]\leq\|G(\theta_{gl}(t))\|^{2}+\sigma^{2}=:M,$
since the gradient updates have bounded variance.
Therefore, we can re-write the bound obtained in the above as
$\mathbb{E}_{t}[\mathcal{L}(\theta_{gl}(t+1))]\leq\mathcal{L}(\theta_{gl}(t))-\eta c\|\nabla\mathcal{L}(\theta_{gl}(t))\|^{2}+L\eta^{2}M/2$
By taking the expectation over the whole iterations, denoted by $\mathbb{E}$, we get
$\mathbb{E}[\mathcal{L}(\theta_{gl}(t+1))]\leq\mathbb{E}[\mathcal{L}(\theta_{gl}(t))]-\eta c\mathbb{E}[\|\nabla\mathcal{L}(\theta_{gl}(t))\|^{2}]+L\eta^{2}M/2$
If we sum over times, from the termalization time $T_{0}$ to the final round T1T-1, and we rearrange the inequalities, we get
$\eta c\sum_{t=T_{0}}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{gl}(t))\|^{2}]\leq\sum_{t=T_{0}}^{T-1}(\mathbb{E}[\mathcal{L}(\theta_{gl}(t))]-\mathbb{E}[\mathcal{L}(\theta_{gl}(t+1))])+(T-T_{0})L\eta^{2}M/2$
Expanding the telescoping sum term on the right-hand-side, it reduces to $\mathbb{E}[\mathcal{L}(\theta_{gl}(T_{0}))]-\mathbb{E}[\mathcal{L}(\theta_{gl}(T)]$. Let $\mathcal{L}^{*}$ be the global optimum of the global loss, then $\mathcal{L}(\theta_{gl}(T)]\geq\mathcal{L}^{*}$ and the expression above becomes
$\eta c\sum_{t=T_{0}}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{gl}(t))\|^{2}]\leq\mathcal{L}(\theta_{gl}(T_{0}))-\mathcal{L}^{*}+(T-T_{0})L\eta^{2}M/2$
Dividing both terms by $\eta c(T-T_{0})$ the inequality becomes
$\dfrac{1}{T-T_{0}}\sum_{t=T_{0}}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{gl}(t))\|^{2}]\leq\dfrac{\mathcal{L}(\theta_{gl}(T_{0}))}{\eta c(T-T_{0})}-\dfrac{\mathcal{L}^{*}}{\eta c(T-T_{0})}+\dfrac{L\eta M}{2c}$
For $T\to\infty$, $T_{0}$ is negligible with respect to $T$. Hence, by finally adopting a change of variable and taking the limit with respect to $T$, we obtain the claim as
$\lim_{T\to\infty}\dfrac{1}{T}\sum_{t=0}^{T}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{gl}(t))\|^{2}]\leq\dfrac{L\eta M}{2c}=:\varepsilon$
Appendix C Implementation Details and Further Experiments